the sum of 2 whole numbers is 9 and the sum of their reciprocals is 1/2. find the numbers
Answer by lwsshak3(6513)

the sum of 2 whole numbers is 9 and the sum of their reciprocals is 1/2. find the numbers
**
let x=one of the numbers
9-x=2nd number
1/x+1/(9-x)=1/2
LCD: 2x(9-x)
18-2x+2x=9x-x^2
x^2-9x+18=0
(x-6)(x-3)=0
x=6
and
x=3
numbers are: 6 and 3